### v1

@@ Abstract
[Abstract]Positivity violations pose a key challenge in the estimation of causal effects, particularly for continuous interventions. Current approaches for addressing this issue include the use of projection functions or modified treatment policies. While effective in many contexts, these methods can result in estimands that potentially do not align well with the original research question, thereby leading to compromises in interpretability. In this paper, we introduce a novel diagnostic tool—the non-overlap ratio—to detect positivity violations. To address these violations while maintaining interpretability, we propose a data-adaptive solution, specially a "most feasible" intervention strategy. Our strategy operates on a unit-specific basis. For a given intervention of interest, we first assess whether the intervention value is feasible for each unit. For units with sufficient support –conditional on confounders– , we adhere to the intervention of interest. However, for units lacking sufficient support –as identified through the assessment of the non-overlap ratio– we do not assign the actual intervention value of interest. Instead, we assign the closest feasible value within the support region. We propose an estimator using g-computation coupled with flexible conditional density estimation to estimate high- and low support regions to estimate this new estimand. Through simulations, we demonstrate that our method effectively reduces bias across various scenarios by addressing positivity violations. Moreover, when positivity violations are absent, the method successfully recovers the standard estimand. We further validate its practical utility using real-world data from the CHAPAS-3 trial, which enrolled HIV-positive children in Zambia and Uganda.
keywords:
data-adaptive interventions; causal inference; positivity violations; HIV treatment
volume: 00footnotetext:
@@ 1 Introduction
Causal inference methods have been extensively studied for binary treatments, where individuals are either exposed to a treatment or assigned to a control condition 1, 2. In this framework, the positivity assumption—requiring that each individual has a non-zero probability of receiving any treatment level under consideration within confounder strata—is crucial for the identifiability and unbiased estimation of causal effects.
@@
Violations of the positivity assumption can be categorized as structural or practical 3, 4, 5. Structural violations occur when an intervention is inherently impossible, whereas practical violations arise when certain treatment levels are theoretically possible but are not observed in the data. Both types lead to non-overlap in the confounder space, where certain confounder patterns have a zero probability, either theoretical or estimated, of receiving specific treatment levels. These violations hinder the estimation of causal effects, making it challenging to draw reliable inferences.
@@
In pharmacoepidemiological and public health studies, researchers often extend their focus to continuous interventions, where the treatment variable can take on a range of values. In this context, the estimation of the causal dose-response curve (CDRC)—which describes the relationship between the continuous treatment and the outcome—has gained increasing attention in the causal inference literature. The CDRC provides a nuanced understanding of causal effects of continuous interventions. However, estimating the CDRC introduces new challenges, particularly because the positivity assumption becomes more stringent with continuous interventions.
@@
For continuous interventions, the positivity assumption is commonly violated due to the infinite dimensionality of the treatment space. This makes it virtually impossible to observe all potential treatment levels across all confounder combinations. The resulting sparsity in certain regions of the intervention space can lead to unstable or biased estimates, a challenge frequently overlooked in existing studies.
@@
While methods adapted from binary treatments can be applied, they have limitations. Trimming, for example, excludes regions of the confounder space with poor overlap 6, removing individuals with extreme probabilities of treatment given their confounders. However, this approach may result in the loss of valuable data.
@@
To tackle these challenges, researchers have developed specialized methods, including projection functions and g-computation-type plug-in estimators, which aim to mitigate the violation of the positivity assumption 4. One such approach is the weighted CDRC, which reweights and redefines the estimand of interest based on functions of the conditional support for the respective interventions 7. Additionally, machine learning techniques have proven effective in enhancing estimation accuracy by minimizing the risk of bias caused by model misspecification 8.
Another notable strategy involves modified treatment policies (MTPs) 9, 10, which redefine the intervention of interest based on the natural intervention (i.e., the level of treatment actually received). While MTPs can sometimes, if designed cleverly, substantially reduce the reliance on the positivity assumption, they (typically) alter the original research question, which may not be ideal in many cases (including the data example discussed in Section 5).
In this paper, we aim to improve both interpretability and applicability in addressing positivity violations for continuous interventions. Our contributions are:
@@
    Defining and diagnosing positivity violations in practice: we introduce a practical framework for diagnosing positivity violations in continuous interventions. Specifically, we define the support using the high-density region of the conditional intervention distribution and propose a diagnostic metric, the non-overlap ratio, to quantify the severity of positivity violations.
@@
    Introducing novel estimands to address positivity violations: we propose a framework of data-adaptive interventions—specifically, the most feasible intervention—to construct a new estimand that replaces unfeasible interventions with the closest feasible values based on the observed data. This approach improves both interpretability and applicability, allowing researchers to approximate the standard CDRC without significantly deviating from the original intervention strategy.
@@ 2.1 Notation
Consider cross-sectional data at a single time point. In this setup:
  • $Y$ is the outcome, with $\mathcal{Y}$ representing the outcome space.
  • $A$ stands for the intervention, with $\mathcal{A}$ being the intervention space.
  • $\mathbf{L}=\{L_{1},L_{2},\dots,L_{q}\}$ represents a set of $q$ confounders, where $q\in\mathbb{N}$. The space of all such confounders is given by $\mathcal{L}$.
We use uppercase letters to denote random variables and lowercase letters for their particular realizations. The temporal order of these variables ranges from $L_{1}$ to $L_{q}$, followed by $A$ and $Y$. Central to our analysis is the concept of counterfactual outcomes. For any subject $i$, the counterfactual outcome $Y^{a}_{i}$ denotes the potential outcome had the intervention $A_{i}$ been set to a specific level $a$ in the intervention space $\mathcal{A}$.
Given a probability space defined by $(\Omega,\mathcal{F},P)$, where $\Omega$ is the product space ×$\mathcal{L}\times\mathcal{A}\times\mathcal{Y}$, the observed data structure can be expressed as:
$O=(\mathbf{L},A,Y)\overset{iid}{\sim}P_{0}$
Here, $O$ is distributed according to $P_{0}$, and the observations are independent and identically distributed.
For a sample of $n$ observations from the distribution $O\sim P_{0}$, the dataset is given by $\mathcal{D}=(o_{1},\dots,o_{n})$.
Further notation:
  • The marginal probability distribution of treatment $A$ is represented by $P_{A}$, and its probability density function is $f_{A}$.
  • For confounders $\mathbf{L}$, their marginal probability distribution and density functions are $P_{\mathbf{L}}$ and $f_{\mathbf{L}}$, respectively.
  • The conditional probability distribution of the treatment given the confounders is $P_{A|\mathbf{L}}$, and its density function is $f_{A|\mathbf{L}}$.
2.2 Consequences of Positivity Violations
In causal inference, assumptions like consistency, exchangeability, and positivity are critical for identifying causal effects. Among these, positivity, often considered as a well-understood assumption, plays a unique role in ensuring that causal treatment effects can be properly identified.
The widely accepted positivity assumption, often referred to as the strong positivity assumption, states that each possible treatment level must have a positive probability of occurring within every stratum of the confounders $\mathbf{L}$ 4:
@@ 2.2 Consequences of Positivity Violations
The positivity assumption requires a non-zero infimum of density in the entire intervention space $\mathcal{A}$, yet depending on the intervention distribution. While much attention has been focused on understanding positivity as an abstract assumption for identification, there is a gap in the literature when it comes to detecting positivity violations in real-world data. Moreover, even when the identification assumptions are met, regions with sparse data can significantly impact the extent of bias and precision at the estimation stage, see Appendix Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies for details. Reliable causal estimation becomes challenging due to inadequate representation across the entire intervention domain for all confounder patterns, which can introduce biases and limit the generalizability of the estimated causal effects.
2.3 Determination of Positivity Violations
Determining whether data sufficiently supports a given intervention requires assessing if the density of the data is so low that it becomes nearly impossible to observe it in finite sample settings. This leads to a crucial insight: introducing a threshold, $c$, for the density function to identify potential positivity violations. Given a specific realization of confounders, $\mathbf{L}=\mathbf{l}$, an intervention of interest $a$ can be classified as follows:
$&\text{supported},&\text{ if }f(a|\mathbf{l})>c,\\ &\text{not supported},&\text{ if }f(a|\mathbf{l})\leqslant c.\end{aligned}\right.$ (2)
@@ 2.3 Determination of Positivity Violations
Establishing a fixed threshold $c$ for continuous interventions thus presents significant challenges. Recognizing this, we propose a novel approach as a key contribution of this paper. This approach redefines and more accurately determines positivity violations in practical settings by utilizing a dynamic diagnostic measure.
2.3.1 Proposal 1: Definition of support and diagnostic using High-Density Region (HDR)
Conditional Support: high-density region
To address the issue of poorly supported intervention areas, we propose a method that dichotomizes the normalized $f$-values. This approach involves identifying regions of the intervention space, $\mathcal{A}$, where the support is either strong or lacking, based on a probability threshold. Specifically, we partition $\mathcal{A}$ into two parts, $\mathcal{A}_{\alpha}(\mathbf{l})$ and its complement $\mathcal{A}^{c}_{\alpha}(\mathbf{l})$, according to a specified threshold depending on $\mathbf{l}$. Given the threshold for detection of non-overlap, named the coverage level, $\alpha$, the region $\mathcal{A}_{\alpha}(\mathbf{l})$ is defined such that the probability measure satisfies $P(\mathcal{A}_{\alpha}(\mathbf{l}))=\alpha$. An intervention $a$ can then be categorized as follows:
$&\text{supported},&\text{ if }a\in\mathcal{A}_{\alpha}(\mathbf{l}),\\ &\text{not supported},&\text{ if }a\notin\mathcal{A}_{\alpha}(\mathbf{l}).\end{aligned}\right.$ (3)
Defining $\mathcal{A}_{\alpha}(\mathbf{l})$ as the high-density region—essentially the practical support—is a relatively straightforward process. This is accomplished by leveraging the concept of high-density regions, proposed by Hyndman 11. Specifically, the high-density region for a continuous intervention is defined by:
$\mathcal{A}_{\alpha}(\mathbf{l})=\left\{a\in\mathcal{A}:f(a|\mathbf{l})\geqslant f_{\alpha}\right\},$ (4)
where $f_{\alpha}$ is the adaptive threshold constant, determined such that the condition $P(\mathcal{A}_{\alpha}(\mathbf{l}))=\alpha$ is satisfied.
@@ 2.3.1 Proposal 1: Definition of support and diagnostic using High-Density Region (HDR)
An illustrative example of this concept is provided in Figure 1, where the green regions represent the conditional support, indicating where a intervention is well-supported by the data. In contrast, the grey regions highlight areas with positivity violations, where a intervention is poorly supported. The total probability of the green region sums to the coverage level $\alpha=95\%$ in this case.
Figure 1: Illustration of using the high-density region as conditional support
Diagnostic measure: non-overlap ratio
We introduce a diagnostic measure called the non-overlap ratio to quantify positivity violations. The non-overlap ratio is defined as a function $\tau:\mathcal{A}\to[0,1]$, given by:

### v2

@@ Abstract
Positivity violations pose a key challenge in the estimation of causal effects, particularly for continuous interventions. Current approaches for addressing this issue include the use of weights or modified treatment policies. While effective in many contexts, these methods can result in estimands that do not always align well with the original research question, thereby compromising interpretability.
In this paper, we introduce a novel diagnostic tool—the non-overlap ratio—to detect positivity violations. To address these violations while maintaining interpretability, we propose a data-adaptive solution, specifically the “most feasible” intervention strategy.
Our strategy operates on a unit-specific basis. For a given intervention of interest, we first assess whether the intervention value is feasible for each unit. For units with sufficient support—conditional on confounders—we adhere to the intervention of interest. However, for units lacking sufficient support, we do not assign the actual intervention value of interest. Instead, we assign the closest feasible value within the support region. The non-overlap ratio provides a diagnostic summary of such support across the population.
We propose an estimator using g-computation coupled with flexible conditional density estimation to identify high- and low-support regions and to estimate this new estimand. Through simulations, we demonstrate that our method effectively reduces bias across various scenarios by addressing positivity violations. Moreover, when positivity violations are absent, the method successfully recovers the standard estimand. We further validate its practical utility using real-world data from the CHAPAS-3 trial, which enrolled HIV-positive children in Zambia and Uganda.
Keywords: data-adaptive interventions; causal inference; positivity violations; HIV treatment
@@ 1 Introduction
Causal inference methods have been extensively studied for binary treatments, where individuals are either exposed to a treatment or assigned to a control condition [21, 17]. In this framework, the positivity assumption—requiring that each individual has a non-zero probability of receiving any treatment level under consideration within confounder strata—is crucial for the identifiability and unbiased estimation of causal effects.
@@
Violations of the positivity assumption can be categorized as structural or practical [38, 27, 37]. Structural violations occur when an intervention is inherently impossible, whereas practical violations arise when certain treatment levels are theoretically possible but are not observed in the data. Both types lead to non-overlap in the confounder space, where certain confounder patterns have a zero probability—either theoretical or estimated—of receiving specific treatment levels. These violations hinder the estimation of causal effects, making it challenging to draw reliable inferences.
@@
In pharmacoepidemiological and public health studies, researchers often extend their focus to continuous interventions, where the treatment variable can take on a range of values. In this context, the estimation of the causal dose–response curve (CDRC)—which describes the relationship between the continuous treatment and the outcome—has gained increasing attention in the causal inference literature. The CDRC provides a nuanced understanding of the causal effects of continuous interventions. However, estimating the CDRC introduces new challenges, particularly because the identification of the CDRC necessitates a more restrictive positivity assumption, requiring that the conditional treatment density is greater than zero—which, when interpreted strictly, is never fully met.
@@
For continuous interventions, the positivity assumption is commonly violated due to the infinite dimensionality of the treatment space. This makes it virtually impossible to observe all potential treatment levels across all confounder combinations. The resulting sparsity in certain regions of the intervention space can lead to unstable or biased estimates—a challenge frequently overlooked in existing studies.
@@
While methods adapted from binary treatments can be applied, they have limitations. Trimming, for example, excludes regions of the confounder space with poor overlap [7], removing individuals with extreme probabilities of treatment given their confounders. However, this approach may result in the loss of valuable data.
@@
To tackle these challenges, researchers have developed specialized methods, including projection functions and g-computation-type plug-in estimators, which aim to mitigate the violation of the positivity assumption [27]. One such approach is the weighted CDRC, which reweights and redefines the estimand of interest based on functions of the conditional support for the respective interventions [31]. Additionally, machine learning techniques have proven effective in enhancing estimation accuracy by minimizing the risk of bias caused by model misspecification [25].
Another notable strategy involves modified treatment policies (MTPs)[11, 9], which specify systematic modifications to the observed (natural) treatment, with the goal of avoiding interventions among individuals for whom the treatment would be infeasible. When designed carefully, MTPs can substantially reduce reliance on the positivity assumption. However, they typically target a different causal estimand than the causal dose–response curve (CDRC), which may or may not align with the scientific objectives of a particular study—including the empirical example discussed in Section 5. MTPs are most useful when the modified intervention itself is of substantive scientific interest. Haneuse and Rotnitzky provide an example of a very well-motivated MTP [11].
While double robustness and semiparametric efficiency are desirable properties, they are not always achievable when estimating the full CDRC under continuous interventions. In fact, under only mild smoothness conditions, the CDRC is not pathwise differentiable [8], precluding the construction of root-$n$ consistent or doubly robust estimators without additional structure. Recent work has developed doubly robust estimators under assumptions such as differentiability [22], monotonicity [36], or smooth invertibility [9], but these assumptions may not be justified in many applied settings—especially when the estimand is defined by the scientific question and not chosen for statistical convenience.
In this paper, we pursue a complementary goal: to preserve the original CDRC as the target estimand while improving its identifiability and estimation by redefining the intervention space in a data-adaptive way. This leads to a natural plug-in strategy that avoids extrapolation into unsupported regions, thereby enhancing both interpretability and applicability in the presence of positivity violations. Although such estimators are not doubly robust, they remain closely aligned with the scientific question, offer clear interpretability, and reduce sensitivity to model misspecification by focusing estimation in well-supported regions. Our specific contributions are:
@@
    Defining and diagnosing positivity violations in practice: we introduce a framework for assessing positivity in continuous interventions. Specifically, we define the support using the highest density region of the conditional intervention distribution and propose the non-overlap ratio—a population-level diagnostic that summarizes the extent of non-overlap across confounder strata, thereby quantifying the severity of positivity violations.
@@
    Introducing novel estimands to address positivity violations: we propose a framework of data-adaptive interventions—specifically, the most feasible intervention—to construct a new estimand that replaces infeasible interventions with the closest feasible values based on the observed data. This approach improves both interpretability and applicability, allowing researchers to approximate the standard CDRC without significantly deviating from the original intervention strategy.
@@ 2.1 Notation
We consider the following setup:
  • $Y$: outcome, taking values in $\mathcal{Y}$.
  • AA: treatment (or exposure), taking values in $\mathcal{A}$.
  • $\mathbf{L}=(L_{1},\dots,L_{q})$: vector of $q\in\mathbb{N}$ confounders, taking values in $\mathcal{L}$.
  • Temporal ordering: $L_{1},\dots,L_{q}\rightarrow A\rightarrow Y$.
Uppercase letters denote random variables, and lowercase letters their realizations. We adopt the potential outcome framework [29]: for subject $i$, $Y_{i}^{a}$ denotes the counterfactual outcome that would be observed if $A_{i}$ were set to level $a\in\mathcal{A}$.
Let $O=(\mathbf{L},A,Y)$ denote a generic observation, with true distribution $P_{0}$. The observed data are
$O_{1},\dots,O_{n}\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;P_{0}.$
We use the following notation for distributions and densities:
  • The marginal distribution of $A$ is $P_{A}$, with density $f_{A}$.
  • The marginal distribution of $\mathbf{L}$ is $P_{\mathbf{L}}$, with density $f_{\mathbf{L}}$.
  • The conditional distribution of $A$ given $\mathbf{L}$ is $P_{A|\mathbf{L}}$, with density $f_{A|\mathbf{L}}$.
Throughout, $P_{0}$ denotes the true distribution, $P$ a generic distribution (in theoretical results), and $P_{n}$ the empirical distribution.
2.2 Determination of Positivity Violations
In causal inference, assumptions such as consistency, exchangeability, and positivity are critical for identifying causal effects. Among these, positivity—often regarded as straightforward—plays a unique role in ensuring that causal treatment effects can be properly identified.
Formally, the positivity assumption requires that every treatment level has positive probability within each stratum of the confounders $\mathbf{L}$ [27]:
$\underset{a\in\mathcal{A}}{\mathrm{inf}}\,f_{A|\mathbf{L}}(a\mid\mathbf{L})>0,\quad\text{a.e.}$ (1)
That is, the conditional treatment density must be bounded away from zero across the entire intervention space $\mathcal{A}$, with the precise requirement depending on the distribution of the intervention. While positivity is typically discussed as an abstract identification assumption, less attention has been paid to detecting violations in practice.
Even when identification holds, sparse data in certain regions of the intervention space can severely affect bias and precision at the estimation stage (see Appendix A for details). Reliable causal estimation becomes difficult when parts of the intervention domain are poorly represented for some confounder patterns, introducing bias and limiting the generalizability of the estimated causal effects.
Determining whether the data provide sufficient support for a given intervention requires assessing whether the density is so low that observing it in finite samples becomes nearly impossible. This motivates introducing a threshold $c$ on the density function to identify potential positivity violations, as proposed by Schomaker et al. [31]. Given a specific realization of confounders, $\mathbf{L}=\mathbf{l}$, an intervention of interest $a$ can be classified as:
@@ 2.2 Determination of Positivity Violations
For binary interventions, (near-)violations occur when the conditional probability of observing an intervention is extremely small or zero. In this discrete case, densities coincide with probabilities, making violations relatively straightforward to detect. For continuous interventions, however, the situation is more complex. The density function must satisfy
$\int_{\mathcal{A}}f(a)\,da=1,$
so the scale of $f$ depends on the size of the intervention space $\mathcal{A}$. This variability complicates the use of static thresholds; for instance, defining interventions as unsupported when $f<0.001$ or as sufficiently supported when $f>0.1$ may not be meaningful across different contexts.
Because of this, fixed thresholds are problematic for continuous interventions. To address this limitation, we propose a novel dynamic diagnostic measure that redefines and more accurately determines positivity violations in practical settings.
2.3 Proposal 1: Definition of Support and Diagnostic Using Highest Density Regions (HDRs)
2.3.1 Conditional support via highest density regions
To address the issue of poorly supported intervention areas, we propose a method that dichotomizes the normalized density values. This approach identifies regions of the intervention space $\mathcal{A}$ where support is either adequate or lacking, based on a probability threshold. Specifically, we partition $\mathcal{A}$ into two sets, $\mathcal{A}_{\alpha}(\mathbf{l})$ and its complement $\mathcal{A}^{c}_{\alpha}(\mathbf{l})$, according to a specified threshold depending on $\mathbf{l}$. Given a threshold for non-overlap detection, referred to as the support level $\alpha$, the set $\mathcal{A}_{\alpha}(\mathbf{l})$ is defined such that
$P(\mathcal{A}_{\alpha}(\mathbf{l}))=\alpha.$
An intervention $a$ can then be categorized as
@@ 2.3.1 Conditional support via highest density regions
We define $\mathcal{A}_{\alpha}(\mathbf{l})$ using highest density regions (HDRs), which represent the practical support of the intervention. This concept, introduced by Hyndman [20], is formally given by
$\mathcal{A}_{\alpha}(\mathbf{l})=\left\{a\in\mathcal{A}:f_{A|\mathbf{L}}(a\mid\mathbf{l})\geqslant f_{\alpha}\right\},$ (4)
where $f_{\alpha}$ is a threshold chosen such that
$P_{A|\mathbf{L}=\mathbf{l}}\!\left(\mathcal{A}_{\alpha}(\mathbf{l})\right)=\alpha.$
@@
Figure 1 illustrates this idea: the green regions indicate the conditional support where interventions are well-supported by the data, while the grey regions highlight areas of positivity violations. In this example, the total probability of the green region corresponds to a support level of $\alpha=0.95$.
Figure 1: Illustration of using highest density regions as conditional support. Green areas indicate well-supported interventions, while grey areas correspond to positivity violations.
2.3.2 Diagnostic measure: the non-overlap ratio
To assess positivity in practice, we introduce the non-overlap ratio. This diagnostic summarizes, at the population level, the extent to which a candidate intervention value $a$ falls outside the conditional support region defined by the HDRs. Formally, we define a function $\tau:\mathcal{A}\to[0,1]$ by
$\tau(a):=\int\mathbbm{1}\{a\notin\mathcal{A}_{\alpha}(\mathbf{L})\}\,dP_{\mathbf{L}}.$ (5)
Here, $\tau(a)$ represents the proportion of individuals whose covariate profiles $\mathbf{L}$ render the intervention value $a$ unsupported, i.e., lying outside the practical support region where the positivity assumption is satisfied.
The non-overlap ratio takes values between 0 and 1:
  • Values near 0 indicate that $a$ is well-supported across the population, with few or no positivity violations at the chosen support level.
  • Values near 1 indicate that $a$ is largely unsupported, reflecting severe positivity violations and suggesting that causal inferences at this value are unreliable.
By aggregating individual-level support into a single summary, the non-overlap ratio provides a population-level diagnostic for evaluating the plausibility of the positivity assumption. Its practical use will be illustrated in the following sections.
3 Estimands to Address Positivity Violations
3.1 Standard Causal Dose–Response Curve
Estimand 1 (Standard) The standard causal dose–response curve is the function $m^{\text{standard}}:\mathcal{A}\to\mathcal{Y}$ defined by $m^{\text{standard}}(a)=\mathbb{E}(Y^{a}),$ which can be identified as $m^{\text{standard}}(a)$ $\overset{(i)}{=}\,\mathbb{E}\!\left[\mathbb{E}(Y^{a}\mid A=a,\mathbf{L})\right]$ (6) $\overset{(ii)}{=}\,\mathbb{E}\!\left[\mathbb{E}(Y\mid A=a,\mathbf{L})\right]$ $=\int\mathbb{E}(Y\mid A=a,\mathbf{L}=\mathbf{l})\,dP_{\mathbf{L}}(\mathbf{l}),$ where equality (i) follows by the law of iterated expectations and conditional exchangeability, and equality (ii) by consistency and the positivity assumption (Equation 1).
The standard causal dose–response curve, reliant on the positivity assumption in Equation 1, faces serious challenges due to frequent violations, particularly with continuous interventions. As the intervention range widens, regions of the confounder space often have limited or no support, so that $\hat{f}(a\mid\mathbf{L})$ approaches zero. In such cases, identification fails and estimation becomes unstable.
This difficulty is magnified in practice because continuous exposures span infinitely many potential values, making it impossible to observe all treatment levels across every confounder pattern in finite samples. Consequently, estimation procedures are forced to extrapolate into sparsely or never observed regions of the intervention space. Such extrapolation produces unstable estimates, inflates standard errors, and introduces bias, thereby compromising both the validity and the generalizability of the estimated dose–response relationship.
3.2 Existing Estimand to Address Positivity Violations:
Weighted Causal Dose–Response Curve
Estimand 2 (Weighted CDRC) Schomaker et al. [31] propose the weighted CDRC, defined as $m^{\text{weighted}}(a)$= $\,\mathbb{E}\!\left(Y^{a}\,\middle|\,f(a\mid\mathbf{L})>c\right)P\!\big(f(a\mid\mathbf{L})>c\big)$ (7) $+\mathbb{E}\!\left(Y^{a}\,w(a\mid\mathbf{L})\,\middle|\,f(a\mid\mathbf{L})\leq c\right)P\!\big(f(a\mid\mathbf{L})\leq c\big),\qquad\forall a\in\mathcal{A},$ which admits the identification formula $m^{\text{weighted}}(a)=\int\mathbb{E}(Y\mid A=a,\mathbf{L}=\mathbf{l})\,w(a,\mathbf{l})\,dP_{\mathbf{L}}(\mathbf{l}),$ (8) where the weight function $w(a,\mathbf{l})$ is given by $w(a,\mathbf{l})=\begin{cases}1,&\text{if }f(a\mid\mathbf{l})>c,\\[6.0pt] \dfrac{f(a\mid\mathbf{l})}{f(a)},&\text{if }f(a\mid\mathbf{l})\leq c.\end{cases}$ (9)
The weighted CDRC mitigates positivity violations by coinciding with the standard estimand in regions with adequate support, and by redefining it in regions with limited support through a projection based on conditional densities. This ensures identifiability and avoids extrapolation into unsupported regions. However, it changes the target parameter: the resulting curve no longer represents the causal effect of assigning all individuals to treatment level $a$, but rather a blend of causal and weighted associational components. As a consequence, the weighted CDRC may diverge from the scientific estimand of interest, and its causal interpretation becomes less transparent.
More broadly, this approach highlights a fundamental trade-off in addressing positivity violations: one can either preserve the original scientific estimand but accept estimation that is potentially unstable, highly variable, or biased in regions with sparse data, or redefine the estimand to guarantee identifiability at the cost of causal interpretability. The weighted CDRC takes the latter path, producing well-defined estimates that avoid extrapolation into unsupported regions, but in doing so it targets a modified scientific question. In applied settings, researchers must therefore balance the benefits of improved precision and robustness against the loss of a direct causal interpretation when deciding whether to adopt the weighted CDRC.
3.3 Novel Estimands to Address Positivity Violations: Causal Dose–Response Curve with Data-Adaptive Intervention
To address these challenges, we propose a novel estimand that mitigates the impact of positivity violations by focusing on well-supported regions of the intervention space, guided by the diagnostic tool proposed in the previous section. This approach relaxes the strict positivity requirement by focusing on empirically supported regions of the intervention space, while maintaining interpretability of the resulting causal estimand.
3.3.1 Data-Adaptive Intervention
Unlike static interventions, which assign a fixed treatment $a$ to all units, we introduce a more refined approach called the data-adaptive intervention. This strategy customizes the intervention for each individual based on their confounders and the observed data.
For a given oracle intervention value $a$, the idea can be summarized schematically as